Messana Corporation reported the following data for the month of August:

Inventories: Beginning Ending
     Raw materials $36,000 $24,000
     Work in process $23,000 $17,000
     Finished goods $37,000 $55,000
Additional information:
      Raw materials purchases $61,000
      Direct labor cost $94,000
      Manufacturing overhead cost incurred $54,000
Manufacturing overhead cost applied to Work in Process $56,000
     

The cost of goods manufactured for August is:

$227,000

$229,000

$219,000

$217,000

Messana Corporation reported the following data for the month of August:

Inventories: Beginning Ending
     Raw materials $36,000 $24,000
     Work in process $23,000 $17,000
     Finished goods $37,000 $55,000
Additional information:
      Raw materials purchases $61,000
      Direct labor cost $94,000
      Manufacturing overhead cost incurred $54,000
    Manufacturing overhead cost applied to Work in Process $56,000
     

The adjusted cost of goods sold that appears on the income statement for August is:

$229,000

$209,000

$211,000

$247,000
Answer #1

Raw materials available = Beg raw material inventory + Raw materials purchases - Ending raw materials inventory

= 36,000+61,000-24,000 = 73,000

Raw materials used in production = 73,000

Total manufacturing Cost = Raw materials used + Direct Labor + Manufacturing Overhead applied

= 73,000 + 94,000 + 56,000

= 223,000

Cost of Goods manufactured = Beginning wip + Total manufacturing costs - Ending wip

= 23,000 + 223,000 - 17,000

= 229,000

Option B is the answer

Unadjusted cost of goods sold

= Beginning finished goods inventory + Cost of Goods manufactured - Ending finished goods inventory

= 37,000 + 229,000 - 55,000

= 211,000

Adjusted cost of goods sold

= Unadjusted cost of goods sold - Over applied overhead

= 211,000 - (56,000-54,000)

= 211,000 - 2,000

= 209,000

Option B is the answer